OpenStudy (shyniah365):

can someone tell me What is the first step in solving the equation Three times p minus three equals two p plus seven ?

OpenStudy (anonymous):

3p-3=2p+7 first is regroup the p term on one side and the constants on the other side 3p-2p=7+3 the simplify to find p (:

OpenStudy (shyniah365):

thank you

OpenStudy (anonymous):

np (:

OpenStudy (shyniah365):

i trying to finish a test

OpenStudy (anonymous):

goodluck! u can do it! :D

OpenStudy (shyniah365):

[k-7=-8 +8 +7 k=0 what is the error in this problem

OpenStudy (anonymous):

k-7=-8 k=-8+7 k=-1 u forgot the equal sign that why it made u confused